OpenStudy (anonymous):

PLEASE HELP I HAVE ONLY TILL 11:59 TO GET THESE QUESTIONS DONE. What is the common difference of a 57-term arithmetic sequence where the first term is -25 and the sum is 17,727? (1 points) Answer 11 11.5 12 12.5 What is the sum of a 14-term arithmetic sequence where the last term is 30 and the common difference is -5? (1 points) Answer 875 812.5 455 422.5

OpenStudy (anonymous):

12

OpenStudy (stacey):

Use the formulas for the sum (S) and last term (l) S = n[2a+(n-1)d]/2 l = a + (n-1)d where a=first term, n=number of terms, d=common difference